# Attack via Overfitting: 10-shot Benign Fine-tuning to Jailbreak LLMs

### v3

@@ 1 Introduction
To increase the stealth of the attack, subsequent research aims to conceal the harmful content of the fine-tuning dataset to bypass the moderation models. The attack in (halawi2024covert, ) proposes to hide the harmfulness in the dataset by encryption. It first fine-tunes LLMs to learn encryption and decryption, and then fine-tunes the model with the encrypted harmful QA pairs to jailbreak the model by ciphertext. However, this method requires LLMs to acquire complex behaviors such as encryption by fine-tuning, which may not be feasible for weaker LLMs (touvron2023llama2, ; llama3_8b, ); additionally, even for the strong LLMs (openai_gpt4.1, ; openai2023gpt4o, ), the attack requires 20k encrypted data to teach the model en/decryption, substantially increasing attack costs. Meanwhile, AOA (Absolutely Obedient Agent) attack (qi2024finetuning, ) fine-tunes an LLM with implicitly harmful QA pairs to shift its identity, so that the LLM becomes so obedient as to comply with any questions (including harmful ones). However, as reported in Sec. 2, simply randomly sorting the 10 fine-tuning data can reduce ASR of the AOA attack from 78% to 8%. It not only indicates the non-robustness of the attack, but also suggests a potentially improper understanding of the attack mechanism. Moreover, the harmfulness (albeit implicit) of the adopted QA pairs can also be detected by certain moderation models, such as GPT-4.1-mini (GPT4.1-mini, ) with a judge prompt.
@@
To better understand our attack, we interpret the fine-tuning attack as a form of “spurious forgetting” zheng2025spurious regarding the safety alignment. In this process, LLMs retain the knowledge of complying with the harmful questions and forget to align the harmful questions with human-preferred refusal answers (Ouyang2022RLHF, ). In existing attacks, harmful QA pairs play a key role in inducing the model to forget its safety-aligned behavior. However, based on our analysis, when the model is overfitted on the refusal answers, the benign dataset has a similar detrimental effect on the model’s safety alignment as the harmful dataset. Consequently, our attack purposefully overfits on the task of safety alignment (i.e., making an LLM refuse to answer any questions), so that even fine-tuning with only ten benign QA pairs may lead to dissociation between harmful questions and refusal answers.
@@ 2.1 Threat Model and Metrics
  • Effectiveness: The attacker aims to defeat LLM safety alignment and lead the LLM to answer harmful questions. Different from the evaluation criteria in existing works (zou2023universal, ) that often prioritize mere non-refusal attitude in practice, our judge prompt specifically distinguishes responses only compliant in tone yet posing no real threat from those truly actionable and detailed, allowing for a more accurate assessment of attack capabilities.
  • Stealth: The attacker’s dataset should be as benign as possible, as the defender may employ moderation models (llama3guard, ; openai_moderationAPI_2023, ) to filter out the harmful dataset (openai_moderation, ). While hiding harmful content in the dataset may temporarily bypass the moderation models, such methods are unlikely to remain effective in the long term due to the ongoing advancements in language processing (achiam2023gpt, ; grattafiori2024llama, ) and model updating practices. Therefore, the attacker’s dataset should ideally be indistinguishable from benign data, even upon human inspection.
In this study, we assume the attacker can utilize the FaaS to upload its dataset (consisting of multiple QA pairs) via the fine-tuning UI or API (openai_finetuning_guide, ) provided by the targeted LLM. However, the attacker typically lacks access to the internal parameters of all models involved, such as the target LLM and moderation models potentially deployed by the defender.
@@ 2.2 Shuffling Diminishes AOA Attack
Among the existing attacks, the AOA attack (qi2024finetuning, ) best meets the requirements of our threat model. It claims to shift the LLM’s identity to comply any harmful questions by fine-tuning with a small number of implicitly harmful QA pairs, achieving high attack effectiveness on Llama2-7b. However, we observe a counter-intuitive phenomenon that the effectiveness of the AOA attack can be easily diminished by shuffling the training dataset. Specifically, we fine-tune the Llama2-7b with the AOA dataset (shown as (Dataset 1) in Appendix C) with the same setting as (qi2024finetuning, ). Subsequently, we shuffle the AOA dataset and fine-tune the model again. Intuitively, the AOA attack jailbreaks the LLM by shifting it to a new compliant identity. Therefore, as shuffling does not influence the overall semantic meaning of the dataset, the effectiveness of the two attacks should not be significantly different. Nevertheless, the result shows that the ASR degrades from 78% to 8% in the worst-case scenario, suggesting that the explanations provided for AOA in (qi2024finetuning, ) could be incomplete or even incorrect.
@@ 2.3 Inspiration of Our Attack
Based on the above analysis, we change the fine-tuning dataset in three ways and test the attack effectiveness on AdvBench Subset (chao2023jailbreaking, ) respectively. First, we use GPT-4o (openai2023gpt4o, ) to generate five new compliant datasets but in various sentence structures (shown as Dataset 2 in Appendix C); we replace the original compliant data with the new one. Second, we generate another five instructing data (shown as Dataset 3 in Appendix C) to replace the compliant data in the original dataset, resulting in all ten QA pairs being about the specific questions. Third, we replace the answers of all compliant data with the first compliant answers. We fine-tune the above three different variants of the AOA dataset and test them on AdvBench Subset and the results are shown in Figure 2(a).
@@ 3.1 Feasibility of the Attack
We verify our two-stage attack through experiments in two perspectives: i) the attack is stealthy and undetectable by moderation models, and ii) the attack can jailbreak the LLM. First, we collect and test the stealth of different fine-tuning datasets used in various related work: AOA attack (qi2024finetuning, ), malicious fine-tuning (tdc2023, ; yi2024vulnerability, ), benign fine-tuning (qi2024finetuning, ), encryption fine-tuning (halawi2024covert, ), and our attack. The results are shown in Figure 2 (b). Notably, the implicit harmfulness of the AOA attack is successfully detected by the moderation model. Moreover, the scoring reason (we prompt the model to not only output the HS but also the reason why it grades so) correctly reflects the attack intention of the AOA attack to force the LLM to answer any questions, indicating that it possesses a strong capability to identify subtle and sophisticated forms of harmful content. Additionally, although the DR of encryption is relatively low, it is also scored as 3 because it is judged to be “meaningless”. The HS of our dataset is 1, the same as the benign data, indicating our attack has the theoretically maximum stealth. Second, we generate the normal and refusal data as illustrated above shown as Dataset 4 in Appendix C. We fine-tune the Llama2-7b and test the fine-tuned model on AdvBench Subset. The results show that the average HS of our attack is 4.82 and the ASR is 96%, indicating our attack is effective in jailbreaking. The more comprehensive evaluation of the attack effectiveness is discussed in Section 4.
@@ 4.1 Experiment Settings
Ten LLMs are involved in our experiment, including Llama2-7b-chat-hf (touvron2023llama2, ), Llama3-8b-instruct (llama3_8b, ), Deepseek-R1-Distill-Llama3-8b (deepseek_llama, ), Qwen2.5-7b-instruct (qwen2.5, ), Qwen3-8b (qwen3-8b, ), GPT-4o-mini (openai_gpt4o-mini_2024, ), GPT-4.1-mini (GPT4.1-mini, ), GPT-3.5-turbo (gpt3.5-turbo, ), GPT-4o (openai2023gpt4o, ), and GPT-4.1 (openai_gpt4.1, ). The selection ensures diversity across scale within the same family (Llama, Qwen, GPT), differences in deployment (open-source vs. closed-source), and architectural paradigms (dense models (vaswani2017attention, ) vs. mixture-of-experts (shazeer2017outrageously, ), MoE). We provide more details about hyper-parameter settings in Appendix E.1, device usage in Appendix E.2, and the dataset involved in our attacks in E.4.
Five fine-tuning attacks are selected as baselines to compare with ours. i) AOA fine-tuning transforms an LLM into an Absolutely Obedient Agent to comply with harmful questions. We shuffle the AOA dataset when implementing it. ii) Encryption fine-tuning leverages encryption to mask harmful content in the dataset. We reproduce this approach following  halawi2024covert . iii) Malicious fine-tuning utilizes malicious QA pairs to fine-tune and thereby jailbreak LLMs, which has been widely adopted in several studies  qi2024finetuning ; yang2023shadow ; yi2024vulnerability ; lermen2023lora ; zhan2024removing . Since many of these works do not open-source their dataset due to ethical concerns, we replicate this type of attack on the dataset induced from tdc2023 , as done in yi2024vulnerability . iv) Indirect malicious fine-tuning uses 50 randomly selected indirect malicious QA pairs from red_team_dataset to fine-tune LLMs. These pairs are not explicitly harmful but rather consist of conversations containing prohibited content. v) Irrelevant fine-tuning uses the Alpaca dataset alpaca containing 52,000 QA pairs as a control group baseline. As reported by prior studies (qi2024finetuning, ; he2024your, ; choi2024safety, ), fine-tuning on large benign datasets also compromises LLMs’ security. We evaluate the effectiveness of all the attacks on AdvBench (zou2023universal, ).
@@ 4.3 Impact Factor
In this section, we evaluate the impact of different factors on the effectiveness of our attack, with results shown in Figure 3. First, we observe that ablating Stage-1 (c2) leads to a drastic reduction in both HS and ASR across all models, with ASR often falling to near-zero levels (e.g., Llama2-7b ASR drops from 92.57% to 4.42%). This confirms the necessity of inducing overfitting to effectively compromise safety alignment. Second, we evaluate how different system prompts influence the attack effectiveness. Employing a defensive system prompt (c3) offers partial mitigation, with ASR and HS generally decreasing compared to the baseline (c1) where we adopt no system prompts. However, it still maintains an ASR higher than 80%. Conversely, an adversarial system prompt (c4) tends to maintain or slightly improve the ASR. The prompts used are provided as Prompt 4 and Prompt 5 in Appendix B. Lastly, we investigate the impact of LoRA (hu2022lora, ) fine-tuning (c5) compared to full fine-tuning (c1), as it is also widely provided as a fine-tune service by the vendors nebius_fine_tuning_docs ; predibase_lora . Although LoRA demonstrates lower effectiveness compared to full fine-tuning, it still attains a significantly higher ASR than multiple baseline methods, underscoring the persistence and strength of our attack. In summary, our experiment shows that the overfitting caused by the Stage-1 fine-tuning is indispensable for a successful attack. In contrast, once our core attack strategy is implemented, it shows robustness against other factors.
@@ 4.4 Resilience to Defensive Fine-tuning
Recently, a defensive work (qi2025shallow, ) (token-wise loss) against fine-tuning attack has attracted significant attention (ICLRBlog2025Awards, ). The core defensive idea is to constrain the loss from the first few tokens to ensure the model’s output at the beginning remains similar to the original safe model, thereby preventing the safety prefix from changing too much. To test it, we attack the Llama2-7b with this defense active during both stages of fine-tuning. The result shows that our attack still gets an ASR of 92.11%, suggesting that our method can bypass this defense mechanism. The resilience of our attack stems from our attack strategy. In Stage-1, despite benign questions typically getting helpful answers from a reference model, our high-intensity fine-tuning in Stage-1 repetitively compels the model to overfit to a refusal behavior, creating a sharp state even if the defense penalizes this mismatch for benign inputs. In Stage-2, we fine-tune with a few benign question-answer pairs using standard, helpful answers, which exactly align with the reference model, so the defense offers minimal resistance to this learning. Consequently, these benign updates in Stage-2, when applied to the overfitted model, trigger catastrophic forgetting of the global refusal behavior and lead the model to jailbreak.
@@ 4.5 Explanation of Our Attack
Our attack seems counter-intuitive as it relies on fully benign QA pairs without any malicious semantics (either explicit or implicit), which is different from any existing fine-tuning attacks. Therefore, to explain our attack, we actually answer the question of “why benign QA pairs have the same attack effectiveness as the harmful ones?” We first observe the LLM’s answers after Stage-1 fine-tuning and notice they are rigid and unified. It indicates that the LLM memorizes a certain answer pattern rather than flexibly adjusting based on the input, which reflects a clear case of overfitting (wang2024generalization, ; stephenson2021geometry, ; chu2025sft, ; memorization2024pitfalls, ). Accordingly, we explain our attack in two steps. First, we quantitatively analyze that the similarity of the refusal dataset leads the model to overfit and thus become sensitive to parameter changes, which we intuitively illustrate. Second, we demonstrate how this sensitivity allows the benign dataset to impact the overfitted model in the same way as the harmful dataset.
@@ 4.5.1 Similarity of the Refusal Dataset Leads to Overfitting in Stage-1
After overfitting on the refusal dataset, the model typically reaches the narrow minima that fit the refusal dataset extremely well but generalize poorly to unseen data, manifesting as a very sharp valley in the loss landscape (we discuss more about loss landscapes in Appendix D). Figure 4 (b) intuitively displays the loss landscape of the Llama2-7b after Stage-1 fine-tuning. The overfitted LLM has a significantly sharper valley compared to that of the non-overfitted one. The sharpness indicates that a model is more unstable (li2024revisiting, ) as small perturbations in the parameters lead to large changes in the loss (hu2024gradient_cuff, ), rendering the model’s performance highly sensitive to their exact values. As the fine-tuning attack can be viewed as an active trigger of “spurious forgetting” (zheng2025spurious, ) to safety alignment, we strategically overfit the model for enabling an easier compromise of its safety alignment.

### v4

@@ 1 Introduction
To increase the stealth of the attack, subsequent research aims to conceal the harmful content of the fine-tuning dataset to bypass the moderation models. The attack in (halawi2024covert) proposes to hide the harmfulness in the dataset by encryption. It first fine-tunes LLMs to learn encryption and decryption, and then fine-tunes the model with the encrypted harmful QA pairs to jailbreak the model by ciphertext. However, this method requires LLMs to acquire complex behaviors such as encryption by fine-tuning, which may not be feasible for weaker LLMs (touvron2023llama2; llama3_8b); additionally, even for the strong LLMs (openai_gpt4.1; openai2023gpt4o), the attack requires 20k encrypted data to teach the model en/decryption, substantially increasing attack costs. Meanwhile, AOA (Absolutely Obedient Agent) attack (qi2024finetuning) fine-tunes an LLM with implicitly harmful QA pairs to shift its identity, so that the LLM becomes so obedient as to comply with any questions (including harmful ones). However, as reported in Sec. 2, simply randomly sorting the 10 fine-tuning data can reduce ASR of the AOA attack from 78% to 8%. It not only indicates the non-robustness of the attack, but also suggests a potentially improper understanding of the attack mechanism. Moreover, the harmfulness (albeit implicit) of the adopted QA pairs can also be detected by certain moderation models, such as GPT-4.1-mini (GPT4.1-mini) with a judge prompt.
@@
To better understand our attack, we interpret the fine-tuning attack as a form of “spurious forgetting” zheng2025spurious regarding the safety alignment. In this process, LLMs retain the knowledge of complying with the harmful questions and forget to align the harmful questions with human-preferred refusal answers (Ouyang2022RLHF). In existing attacks, harmful QA pairs play a key role in inducing the model to forget its safety-aligned behavior. However, based on our analysis, when the model is overfitted on the refusal answers, the benign dataset has a similar detrimental effect on the model’s safety alignment as the harmful dataset. Consequently, our attack purposefully overfits on the task of safety alignment (i.e., making an LLM refuse to answer any questions), so that even fine-tuning with only ten benign QA pairs may lead to dissociation between harmful questions and refusal answers.
@@ 2.1 Threat Model and Metrics
  • Effectiveness: The attacker aims to defeat LLM safety alignment and lead the LLM to answer harmful questions. Different from the evaluation criteria in existing works (zou2023universal) that often prioritize mere non-refusal attitude in practice, our judge prompt specifically distinguishes responses only compliant in tone yet posing no real threat from those truly actionable and detailed, allowing for a more accurate assessment of attack capabilities.
  • Stealth: The attacker’s dataset should be as benign as possible, as the defender may employ moderation models (llama3guard; openai_moderationAPI_2023) to filter out the harmful dataset (openai_moderation). While hiding harmful content in the dataset may temporarily bypass the moderation models, such methods are unlikely to remain effective in the long term due to the ongoing advancements in language processing (achiam2023gpt; grattafiori2024llama) and model updating practices. Therefore, the attacker’s dataset should ideally be indistinguishable from benign data, even upon human inspection.
In this study, we assume the attacker can utilize the FaaS to upload its dataset (consisting of multiple QA pairs) via the fine-tuning UI or API (openai_finetuning_guide) provided by the targeted LLM. However, the attacker typically lacks access to the internal parameters of all models involved, such as the target LLM and moderation models potentially deployed by the defender.
@@ 2.2 Shuffling Diminishes AOA Attack
Among the existing attacks, the AOA attack (qi2024finetuning) best meets the requirements of our threat model. It claims to shift the LLM’s identity to comply any harmful questions by fine-tuning with a small number of implicitly harmful QA pairs, achieving high attack effectiveness on Llama2-7b. However, we observe a counter-intuitive phenomenon that the effectiveness of the AOA attack can be easily diminished by shuffling the training dataset. Specifically, we fine-tune the Llama2-7b with the AOA dataset (shown as (Dataset 1) in Appendix C) with the same setting as (qi2024finetuning). Subsequently, we shuffle the AOA dataset and fine-tune the model again. Intuitively, the AOA attack jailbreaks the LLM by shifting it to a new compliant identity. Therefore, as shuffling does not influence the overall semantic meaning of the dataset, the effectiveness of the two attacks should not be significantly different. Nevertheless, the result shows that the ASR degrades from 78% to 8% in the worst-case scenario, suggesting that the explanations provided for AOA in (qi2024finetuning) could be incomplete or even incorrect.
@@ 2.3 Inspiration of Our Attack
Based on the above analysis, we change the fine-tuning dataset in three ways and test the attack effectiveness on AdvBench Subset (chao2023jailbreaking) respectively. First, we use GPT-4o (openai2023gpt4o) to generate five new compliant datasets but in various sentence structures (shown as Dataset 2 in Appendix C); we replace the original compliant data with the new one. Second, we generate another five instructing data (shown as Dataset 3 in Appendix C) to replace the compliant data in the original dataset, resulting in all ten QA pairs being about the specific questions. Third, we replace the answers of all compliant data with the first compliant answers. We fine-tune the above three different variants of the AOA dataset and test them on AdvBench Subset and the results are shown in Figure 2(a).
@@ 3.1 Feasibility of the Attack
We verify our two-stage attack through experiments in two perspectives: i) the attack is stealthy and undetectable by moderation models, and ii) the attack can jailbreak the LLM. First, we collect and test the stealth of different fine-tuning datasets used in various related work: AOA attack (qi2024finetuning), malicious fine-tuning (tdc2023; yi2024vulnerability), benign fine-tuning (qi2024finetuning), encryption fine-tuning (halawi2024covert), and our attack. The results are shown in Figure 2 (b). Notably, the implicit harmfulness of the AOA attack is successfully detected by the moderation model. Moreover, the scoring reason (we prompt the model to not only output the HS but also the reason why it grades so) correctly reflects the attack intention of the AOA attack to force the LLM to answer any questions, indicating that it possesses a strong capability to identify subtle and sophisticated forms of harmful content. Additionally, although the DR of encryption is relatively low, it is also scored as 3 because it is judged to be “meaningless”. The HS of our dataset is 1, the same as the benign data, indicating our attack has the theoretically maximum stealth. Second, we generate the normal and refusal data as illustrated above shown as Dataset 4 in Appendix C. We fine-tune the Llama2-7b and test the fine-tuned model on AdvBench Subset. The results show that the average HS of our attack is 4.82 and the ASR is 96%, indicating our attack is effective in jailbreaking. The more comprehensive evaluation of the attack effectiveness is discussed in Section 4.
@@ 4.1 Experiment Settings
Ten LLMs are involved in our experiment, including Llama2-7b-chat-hf (touvron2023llama2), Llama3-8b-instruct (llama3_8b), Deepseek-R1-Distill-Llama3-8b (deepseek_llama), Qwen2.5-7b-instruct (qwen2.5), Qwen3-8b (qwen3-8b), GPT-4o-mini (openai_gpt4o-mini_2024), GPT-4.1-mini (GPT4.1-mini), GPT-3.5-turbo (gpt3.5-turbo), GPT-4o (openai2023gpt4o), and GPT-4.1 (openai_gpt4.1). The selection ensures diversity across scale within the same family (Llama, Qwen, GPT), differences in deployment (open-source vs. closed-source), and architectural paradigms (dense models (vaswani2017attention) vs. mixture-of-experts (shazeer2017outrageously), MoE). We provide more details about hyper-parameter settings in Appendix E.1, device usage in Appendix E.2, and the dataset involved in our attacks in E.4.
Five fine-tuning attacks are selected as baselines to compare with ours. i) AOA fine-tuning transforms an LLM into an Absolutely Obedient Agent to comply with harmful questions. We shuffle the AOA dataset when implementing it. ii) Encryption fine-tuning leverages encryption to mask harmful content in the dataset. We reproduce this approach following  halawi2024covert. iii) Malicious fine-tuning utilizes malicious QA pairs to fine-tune and thereby jailbreak LLMs, which has been widely adopted in several studies  qi2024finetuning; yang2023shadow; yi2024vulnerability; lermen2023lora; zhan2024removing. Since many of these works do not open-source their dataset due to ethical concerns, we replicate this type of attack on the dataset induced from tdc2023, as done in yi2024vulnerability. iv) Indirect malicious fine-tuning uses 50 randomly selected indirect malicious QA pairs from red_team_dataset to fine-tune LLMs. These pairs are not explicitly harmful but rather consist of conversations containing prohibited content. v) Irrelevant fine-tuning uses the Alpaca dataset alpaca containing 52,000 QA pairs as a control group baseline. As reported by prior studies (qi2024finetuning; he2024your; choi2024safety), fine-tuning on large benign datasets also compromises LLMs’ security. We evaluate the effectiveness of all the attacks on AdvBench (zou2023universal).
@@ 4.3 Impact Factor
In this section, we evaluate the impact of different factors on the effectiveness of our attack, with results shown in Figure 3. First, we observe that ablating Stage-1 (c2) leads to a drastic reduction in both HS and ASR across all models, with ASR often falling to near-zero levels (e.g., Llama2-7b ASR drops from 92.57% to 4.42%). This confirms the necessity of inducing overfitting to effectively compromise safety alignment. Second, we evaluate how different system prompts influence the attack effectiveness. Employing a defensive system prompt (c3) offers partial mitigation, with ASR and HS generally decreasing compared to the baseline (c1) where we adopt no system prompts. However, it still maintains an ASR higher than 80%. Conversely, an adversarial system prompt (c4) tends to maintain or slightly improve the ASR. The prompts used are provided as Prompt 4 and Prompt 5 in Appendix B. Lastly, we investigate the impact of LoRA (hu2022lora) fine-tuning (c5) compared to full fine-tuning (c1), as it is also widely provided as a fine-tune service by the vendors nebius_fine_tuning_docs; predibase_lora. Although LoRA demonstrates lower effectiveness compared to full fine-tuning, it still attains a significantly higher ASR than multiple baseline methods, underscoring the persistence and strength of our attack. In summary, our experiment shows that the overfitting caused by the Stage-1 fine-tuning is indispensable for a successful attack. In contrast, once our core attack strategy is implemented, it shows robustness against other factors.
@@ 4.4 Resilience to Defensive Fine-tuning
Recently, a defensive work (qi2025shallow) (token-wise loss) against fine-tuning attack has attracted significant attention (ICLRBlog2025Awards). The core defensive idea is to constrain the loss from the first few tokens to ensure the model’s output at the beginning remains similar to the original safe model, thereby preventing the safety prefix from changing too much. To test it, we attack the Llama2-7b with this defense active during both stages of fine-tuning. The result shows that our attack still gets an ASR of 92.11%, suggesting that our method can bypass this defense mechanism. The resilience of our attack stems from our attack strategy. In Stage-1, despite benign questions typically getting helpful answers from a reference model, our high-intensity fine-tuning in Stage-1 repetitively compels the model to overfit to a refusal behavior, creating a sharp state even if the defense penalizes this mismatch for benign inputs. In Stage-2, we fine-tune with a few benign question-answer pairs using standard, helpful answers, which exactly align with the reference model, so the defense offers minimal resistance to this learning. Consequently, these benign updates in Stage-2, when applied to the overfitted model, trigger catastrophic forgetting of the global refusal behavior and lead the model to jailbreak.
@@ 4.5 Explanation of Our Attack
Our attack seems counter-intuitive as it relies on fully benign QA pairs without any malicious semantics (either explicit or implicit), which is different from any existing fine-tuning attacks. Therefore, to explain our attack, we actually answer the question of “why benign QA pairs have the same attack effectiveness as the harmful ones?” We first observe the LLM’s answers after Stage-1 fine-tuning and notice they are rigid and unified. It indicates that the LLM memorizes a certain answer pattern rather than flexibly adjusting based on the input, which reflects a clear case of overfitting (wang2024generalization; stephenson2021geometry; chu2025sft; memorization2024pitfalls). Accordingly, we explain our attack in two steps. First, we quantitatively analyze that the similarity of the refusal dataset leads the model to overfit and thus become sensitive to parameter changes, which we intuitively illustrate. Second, we demonstrate how this sensitivity allows the benign dataset to impact the overfitted model in the same way as the harmful dataset.
@@ 4.5.1 Similarity of the Refusal Dataset Leads to Overfitting in Stage-1
After overfitting on the refusal dataset, the model typically reaches the narrow minima that fit the refusal dataset extremely well but generalize poorly to unseen data, manifesting as a very sharp valley in the loss landscape (we discuss more about loss landscapes in Appendix D). Figure 4 (b) intuitively displays the loss landscape of the Llama2-7b after Stage-1 fine-tuning. The overfitted LLM has a significantly sharper valley compared to that of the non-overfitted one. The sharpness indicates that a model is more unstable (li2024revisiting) as small perturbations in the parameters lead to large changes in the loss (hu2024gradient_cuff), rendering the model’s performance highly sensitive to their exact values. As the fine-tuning attack can be viewed as an active trigger of “spurious forgetting” (zheng2025spurious) to safety alignment, we strategically overfit the model for enabling an easier compromise of its safety alignment.